# Landau.bib

### v1

@@ 2.4 Entropy dissipation and Fisher information
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $=$ ${}\Big{(}\int_{\mathbb{R}^{3}\times\{|v_{*}|\leq\frac{|v|}{2}\}}+\int_{\mathbb{R}^{3}\times\{|v_{*}|\geq\frac{|v|}{2}\}}\Big{)}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}\Big{(}\frac{|v|}{2}\Big{)}^{\alpha}\int_{\mathbb{R}^{3}\times\{|v_{*}|\leq\frac{|v|}{2}\}}f(x_{*},v_{*})\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $+\Big{(}\frac{|v|}{2}\Big{)}^{\alpha}\int_{\mathbb{R}^{3}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big{(}\frac{|v-v_{*}|}{|v_{*}|}\Big{)}^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}2^{-\alpha}\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+2^{-\alpha}\int_{\mathbb{R}^{3}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big{(}\frac{|v-v_{*}|}{|v_{*}|}\Big{)}^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}.$
    Note that to obtain the last inequality above, we have used the fact that $|v|\geq 1$ and $\alpha<0$. Similar to (2.30), on the domain $\mathbb{R}^{3}\times\{|v_{*}|\geq\frac{|v|}{2}\}$, we have
    $\int_{\mathbb{R}^{3}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big{(}\frac{|v-v_{*}|}{|v_{*}|}\Big{)}^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq\Big{(}\int_{\mathbb{R}^{3}\times\{|v-v_{*}|\geq 1\}}+\int_{\mathbb{R}^{3}\times\{|v-v_{*}|\leq 1\}}\Big{)}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta+|\alpha|}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+\big{(}\int_{\mathbb{R}^{3}}f(x_{*},\cdot)\langle\cdot\rangle^{\delta+|\alpha|}{\,\rm d}x_{*}\big{)}*_{v}(\chi_{B_{1}}|\cdot|^{\alpha})$
    $\leq\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{p}_{v}L^{1}_{x}}\||v|^{-\gamma}\|_{L^{p^{\prime}}(B_{1})}.$
@@
    $\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\lesssim$ ${}\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{p}_{v}L^{1}_{x}}.$
@@
Let $\alpha\in(-3,0)$ and $\delta\in\mathbb{R}$. Let $f\in\mathcal{P}\cap\mathcal{S}(\mathbb{R}^{3}\times\mathbb{R}^{3})$. Then we have
$\sup_{v\in\mathbb{R}^{3}}\|f*_{v}|v-\cdot|^{\alpha}\langle\cdot\rangle^{\delta}\|_{L^{1}(\mathbb{R}^{3})}\lesssim\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})}+\|\langle v\rangle^{\delta}f\|_{L^{p}_{v}L^{1}_{x}(\mathbb{R}^{3}\times\mathbb{R}^{3})}$
for some $p>\frac{3}{3-|\alpha|}$.
@@ 3 Variational characterisation
For simplification, we write the domain as $\mathbb{R}^{3}\times\mathbb{R}^{3}$, and unless otherwise stated, it refers to $\mathbb{T}^{3}\times\mathbb{R}^{3}$ and $\mathbb{R}^{3}\times\mathbb{R}^{3}$.
@@ 3.1 Transport grazing rate equation
We consider curves $(\mu_{t})_{t\in[0,T]}\subset\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ derived by transport and grazing mechanism via a transport grazing rate continuity equation
$\partial_{t}\mu_{t}+v\cdot\nabla_{x}\mu_{t}+\frac{1}{2}\tilde{\nabla}\cdot\mathcal{U}_{t}=0,$ (3.1)
where $\mathcal{U}_{t}\in\mathcal{M}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$ denotes the grazing rate, and $\mathcal{M}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$ denotes the space of vector-valued signed Borel measures with bounded total variation on $\mathbb{R}^{12}$. The space $\mathcal{M}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$ is endowed with the weak* topology induced by $C_{0}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$ with continuous functions vanishing at infinity. We recall again that the gradient operator, $\tilde{\nabla}$, and the associated divergence operator, $\tilde{\nabla}\cdot$, are defined, respectively, in (1.11) and (1.12). Also, we have denoted
@@
the Lebesgue measure on $(\mathbb{R}^{3}\times\mathbb{R}^{3})^{2}$.
@@
    $\mu_{t}:[0,T]\to\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ is weakly continuous;
@@
    $(\mathcal{U}_{t})_{t\in[0,T]}$ is a family of Borel measures in $\mathcal{M}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$;
@@
    $\int_{0}^{T}{\,\rm d}|\mathcal{U}_{t}|((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2})<\infty$;
@@
    For any $\varphi\in C^{\infty}_{c}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ the following equality holds
    $\frac{d}{dt}\int_{\mathbb{R}^{6}}\varphi{\,\rm d}\mu_{t}-\int_{\mathbb{R}^{6}}v\cdot\nabla\varphi{\,\rm d}\mu_{t}=\frac{1}{2}\int_{\mathbb{R}^{12}}\tilde{\nabla}\varphi{\,\rm d}\mathcal{U}_{t}.$
@@
If the pair $(\mu_{t},\mathcal{U}_{t})$ has densities $f_{t}$ and $U_{t}$ with respect to Lebesgue measure on $\mathbb{R}^{3}\times\mathbb{R}^{3}$ and $(\mathbb{R}^{3}\times\mathbb{R}^{3})^{2}$, we will also write $(f_{t},U_{t})\in\mathcal{TGRE}_{T}$.
We define the function $\alpha:\mathbb{R}_{+}\times\mathbb{R}^{3}\to\mathbb{R}$ by letting
@@
Next we define the action functional for $\mu\in\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ and $\mathcal{U}\in\mathcal{M}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$.
@@
$\mathcal{A}(\mu,\mathcal{U})=\int_{\mathbb{R}^{12}}\alpha\Big{(}\frac{{\,\rm d}\mu_{1}}{{\,\rm d}\lambda},\frac{{\,\rm d}\mathcal{U}}{{\,\rm d}\lambda}\Big{)}{\,\rm d}\lambda,$ (3.3)
@@
Notice that $\mu_{1}\in\mathcal{P}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2})$ and $\mathcal{A}(\mu,\mathcal{U})$ is well-defined since $\alpha$ is 1-homogeneous.
@@
If $\mu=f\mathcal{L}\in\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ and $\mathcal{U}\in\mathcal{M}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$ and $\mathcal{A}(\mu,\mathcal{U})<+\infty$, then there exists a Borel measure $M:(\mathbb{R}^{3}\times\mathbb{R}^{3})^{2}\to\mathbb{R}^{3}$ such that $\mathcal{U}=Mff_{*}\kappa{\,\rm d}\eta=U{\,\rm d}\eta$ and
$\mathcal{A}(\mu,\mathcal{U})=\frac{1}{2}\int_{\mathbb{R}^{12}}|M|^{2}ff_{*}\kappa{\,\rm d}\eta=\frac{1}{2}\int_{\mathbb{R}^{12}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta.$
@@
The proof straightforwardly follows [erbar2023gradient] and [carrillo2024landau]. Let ${\,\rm d}\tilde{\lambda}=\kappa{\,\rm d}\lambda$, ${\,\rm d}\tilde{\mu}_{1}=\kappa{\,\rm d}\mu_{1}$ and ${\,\rm d}\tilde{\mathcal{U}}=\kappa{\,\rm d}\mathcal{U}$. Then 1-homogeneity of $\alpha$ ensures that
$\mathcal{A}(\mu_{t},\mathcal{U}_{t})=\int_{\mathbb{R}^{12}}\alpha\Big{(}\frac{{\,\rm d}\tilde{\mu}_{1}}{{\,\rm d}\tilde{\lambda}},\frac{{\,\rm d}\tilde{\mathcal{U}}}{{\,\rm d}\tilde{\lambda}}\Big{)}{\,\rm d}\tilde{\lambda}<+\infty.$ (3.4)
It is sufficient to show that $\tilde{\mathcal{U}}$ is absolutely continuous with respect to $\tilde{\mu}_{1}$. Let $S\subset\mathbb{R}^{12}$ such that $\tilde{\mu}_{1}(S)=0$, then $\frac{{\,\rm d}\tilde{\mu}_{1}}{{\,\rm d}\tilde{\lambda}}=0$ $\tilde{\lambda}$-almost everywhere. By definition of $\alpha$ in (3.2) and the finiteness (3.4), we have $\frac{{\,\rm d}\tilde{\mathcal{U}}}{{\,\rm d}\tilde{\lambda}}=0$ $\tilde{\lambda}$-almost everywhere. Hence, $\mathcal{U}$ is absolutely continuous with respect to $\tilde{\mu}_{1}$ and $\sigma$. ∎
@@
Let $\mu_{n}\rightharpoonup\mu$ in $\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ and $\mathcal{U}_{n}\overset{*}{\rightharpoonup}\mathcal{U}$ in $\mathcal{M}((\mathbb{R}^{3}\times\mathbb{R}^{3})^{2};\mathbb{R}^{3})$. We have
@@
then for $\gamma\in(-3,1]$, we have
$\int_{0}^{T}\int_{\mathbb{R}^{12}}|U_{t}|\big{(}\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle v_{*}\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle+\langle x_{*}\rangle\big{)}{\,\rm d}\eta{\,\rm d}t\leq\sqrt{C_{A}C_{E}}.$ (3.5)
When $\gamma\in(-3,-2)$, if in addition $f\in L^{1}([0,T];L^{p}_{v}L^{1}_{x})$ for some $p>\frac{3}{3-|2+\gamma|}$, then we have
$\int_{0}^{T}\int_{\mathbb{R}^{12}}|U_{t}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}t\leq\sqrt{CC_{A}},$ (3.6)
where $C\lesssim\|\langle v\rangle^{|2+\gamma|}f\|_{L^{1}([0,T]\times\mathbb{R}^{3}\times\mathbb{R}^{3})}+\|\langle v\rangle^{|2+\gamma|}f\|_{L^{1}([0,T];L^{p}_{v}L^{x})}$.
@@
We first show (3.5). By Cauchy–Schwartz inequality, we have
$\int_{0}^{T}\int_{\mathbb{R}^{12}}|U_{t}|\big{(}\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle\big{)}{\,\rm d}\eta{\,\rm d}t$ (3.7)
$\leq\Big{(}\int_{0}^{T}\int_{\mathbb{R}^{12}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta{\,\rm d}t\Big{)}^{\frac{1}{2}}\times$
$\quad\Big{(}\int_{0}^{T}\int_{\mathbb{R}^{12}}\big{(}\langle v\rangle^{2+|\gamma|}+\langle x\rangle^{2}\big{)}{ff_{*}}\kappa{\,\rm d}\eta{\,\rm d}t\Big{)}^{\frac{1}{2}}$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{12}}|U_{t}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}r$
$\leq{C_{A}}^{\frac{1}{2}}\Big{(}\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa{ff_{*}}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t\Big{)}^{\frac{1}{2}}.$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{12}}\kappa{ff_{*}}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t\lesssim 1+\|f\|_{L^{1}([0,T];L^{p}_{v}L^{1}_{x})}.$

### v2

@@ 2.4 Entropy dissipation and Fisher information
    $\int_{\mathbb{R}^{2d}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $=$ ${}\Big(\int_{\mathbb{R}^{d}\times\{|v_{*}|\leq\frac{|v|}{2}\}}+\int_{\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}}\Big)f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}\Big(\frac{|v|}{2}\Big)^{\alpha}\int_{\mathbb{R}^{d}\times\{|v_{*}|\leq\frac{|v|}{2}\}}f(x_{*},v_{*})\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $+\Big(\frac{|v|}{2}\Big)^{\alpha}\int_{\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big(\frac{|v-v_{*}|}{|v_{*}|}\Big)^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq$ ${}2^{-\alpha}\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{2d})}+2^{-\alpha}\int_{\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big(\frac{|v-v_{*}|}{|v_{*}|}\Big)^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}.$
    Note that to obtain the last inequality above, we have used the fact that $|v|\geq 1$ and $\alpha<0$. Similar to (2.27), on the domain $\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}$, we have
    $\int_{\mathbb{R}^{d}\times\{|v_{*}|\geq\frac{|v|}{2}\}}f(x_{*},v_{*})\Big(\frac{|v-v_{*}|}{|v_{*}|}\Big)^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq\Big(\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\geq 1\}}+\int_{\mathbb{R}^{d}\times\{|v-v_{*}|\leq 1\}}\Big)f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta+|\alpha|}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\leq\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{2d})}+\big(\int_{\mathbb{R}^{d}}f(x_{*},\cdot)\langle\cdot\rangle^{\delta+|\alpha|}{\,\rm d}x_{*}\big)*_{v}(\chi_{B_{1}}|\cdot|^{\alpha})$
    $\leq\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{2d})}+\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{p}_{v}L^{1}_{x}}\||v|^{-\gamma}\|_{L^{p^{\prime}}(B_{1})}.$
@@
    $\int_{\mathbb{R}^{2d}}f(x_{*},v_{*})|v-v_{*}|^{\alpha}\langle v_{*}\rangle^{\delta}{\,\rm d}x_{*}{\,\rm d}v_{*}$
    $\lesssim$ ${}\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{1}(\mathbb{R}^{2d})}+\|\langle v\rangle^{\delta+|\alpha|}f\|_{L^{p}_{v}L^{1}_{x}}.$
@@
Let $\alpha\in(-d,0)$ and $\delta\in\mathbb{R}$. Let $f\in\mathcal{P}\cap\mathcal{S}(\mathbb{R}^{2d})$. Then we have
$\sup_{v\in\mathbb{R}^{d}}\|\langle v\rangle^{\delta}f*_{v}|v|^{\alpha}\|_{L^{1}_{x}}\lesssim\|\langle v\rangle^{\delta}f\|_{L^{1}(\mathbb{R}^{2d})}+\min(\|\langle v\rangle^{\delta}f\|_{L^{1}_{x}L^{p}_{v}},\|\langle v\rangle^{\delta}f\|_{L^{p}_{v}L^{1}_{x}})$
for some $p>\frac{d}{d-|\alpha|}$.
@@ 3 Variational characterisation
For simplification, we write the domain as $\mathbb{R}^{2d}$, and unless otherwise stated, it refers to $\mathbb{T}^{d}\times\mathbb{R}^{d}$ and $\mathbb{R}^{2d}$.
@@ 3.1 Transport grazing rate equation
We consider curves $(\mu_{t})_{t\in[0,T]}\subset\mathcal{P}(\mathbb{R}^{2d})$ derived by transport and grazing mechanism via a transport grazing rate continuity equation
$\partial_{t}\mu_{t}+v\cdot\nabla_{x}\mu_{t}+\frac{1}{2}\widetilde{\nabla}\cdot\mathcal{U}_{t}=0,$ (3.1)
where $\mathcal{U}_{t}\in\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$ denotes the grazing rate, and $\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$ denotes the space of vector-valued signed Borel measures with bounded total variation on $\mathbb{R}^{4d}$. The space $\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$ is endowed with the weak* topology induced by $C_{0}(\mathbb{R}^{4d};\mathbb{R}^{d})$ with continuous functions vanishing at infinity. We recall again that the gradient operator, $\widetilde{\nabla}$, and the associated divergence operator, $\widetilde{\nabla}\cdot$, are defined, respectively, in (1.11) and (1.12). Also, we have denoted
@@
the Lebesgue measure on $\mathbb{R}^{4d}$.
@@
    $\mu_{t}:[0,T]\to\mathcal{P}(\mathbb{R}^{2d})$ is weakly continuous;
@@
    $(\mathcal{U}_{t})_{t\in[0,T]}$ is a family of Borel measures in $\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$;
@@
    $\int_{0}^{T}{\,\rm d}|\mathcal{U}_{t}|(\mathbb{R}^{4d})<\infty$;
@@
    For any $\varphi\in C^{\infty}_{c}(\mathbb{R}^{2d})$ the following equality holds
    $\frac{d}{dt}\int_{\mathbb{R}^{2d}}\varphi{\,\rm d}\mu_{t}-\int_{\mathbb{R}^{2d}}v\cdot\nabla\varphi{\,\rm d}\mu_{t}=\frac{1}{2}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\varphi{\,\rm d}\mathcal{U}_{t}.$
@@
If the pair $(\mu_{t},\mathcal{U}_{t})$ has densities $f_{t}$ and $U_{t}$ with respect to Lebesgue measure on $\mathbb{R}^{2d}$ and $\mathbb{R}^{4d}$, we will also write $(f_{t},U_{t})\in\mathcal{TGRE}_{T}$.
We define the function $\alpha:\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}$ by letting
@@
Next we define the action functional for $\mu\in\mathcal{P}(\mathbb{R}^{2d})$ and $\mathcal{U}\in\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$.
@@
$\mathcal{A}(\mu,\mathcal{U})=\int_{\mathbb{R}^{4d}}\alpha\Big(\frac{{\,\rm d}\mu_{1}}{{\,\rm d}\lambda},\frac{{\,\rm d}\mathcal{U}}{{\,\rm d}\lambda}\Big){\,\rm d}\lambda,$ (3.3)
@@
Notice that $\mu_{1}\in\mathcal{P}(\mathbb{R}^{4d})$ and $\mathcal{A}(\mu,\mathcal{U})$ is well-defined since $\alpha$ is 1-homogeneous.
@@
If $\mu=f\mathcal{L}\in\mathcal{P}(\mathbb{R}^{2d})$ and $\mathcal{U}\in\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$ and $\mathcal{A}(\mu,\mathcal{U})<+\infty$, then there exists a Borel measure $M:\mathbb{R}^{4d}\to\mathbb{R}^{d}$ such that $\mathcal{U}=Mff_{*}\kappa{\,\rm d}\eta=U{\,\rm d}\eta$ and
$\mathcal{A}(\mu,\mathcal{U})=\frac{1}{2}\int_{\mathbb{R}^{4d}}|M|^{2}ff_{*}\kappa{\,\rm d}\eta=\frac{1}{2}\int_{\mathbb{R}^{4d}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta.$
@@
The proof straightforwardly follows [Erb23] and [CDDW24]. Let ${\,\rm d}\widetilde{\lambda}=\kappa{\,\rm d}\lambda$, ${\,\rm d}\widetilde{\mu}_{1}=\kappa{\,\rm d}\mu_{1}$ and ${\,\rm d}\widetilde{\mathcal{U}}=\kappa{\,\rm d}\mathcal{U}$. Then 1-homogeneity of $\alpha$ ensures that
$\mathcal{A}(\mu_{t},\mathcal{U}_{t})=\int_{\mathbb{R}^{4d}}\alpha\Big(\frac{{\,\rm d}\widetilde{\mu}_{1}}{{\,\rm d}\widetilde{\lambda}},\frac{{\,\rm d}\widetilde{\mathcal{U}}}{{\,\rm d}\widetilde{\lambda}}\Big){\,\rm d}\widetilde{\lambda}$$\infty$. (3.4)
It is sufficient to show that $\widetilde{\mathcal{U}}$ is absolutely continuous with respect to $\widetilde{\mu}_{1}$. Let $S\subset\mathbb{R}^{4d}$ such that $\widetilde{\mu}_{1}(S)=0$, then $\frac{{\,\rm d}\widetilde{\mu}_{1}}{{\,\rm d}\widetilde{\lambda}}=0$ $\widetilde{\lambda}-$almost everywhere. By definition of $\alpha$ in (3.2) and the finiteness (3.4), we have $\frac{{\,\rm d}\widetilde{\mathcal{U}}}{{\,\rm d}\widetilde{\lambda}}=0$ $\widetilde{\lambda}-$almost everywhere. Hence, $\mathcal{U}$ is absolutely continuous with respect to $\widetilde{\mu}_{1}$ and $\sigma$. ∎
@@
Let $\mu_{n}\rightharpoonup\mu$ in $\mathcal{P}(\mathbb{R}^{2d})$ and $\mathcal{U}_{n}\overset{*}{\rightharpoonup}\mathcal{U}$ in $\mathcal{M}(\mathbb{R}^{4d};\mathbb{R}^{d})$. We have
@@
then for $\gamma\in(-\gamma_{d},1]$, we have
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U_{t}|\big(\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle v_{*}\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle+\langle x_{*}\rangle\big){\,\rm d}\eta{\,\rm d}t\leq\sqrt{C_{A}C_{E}}.$ (3.5)
When $\gamma\in(-\gamma_{d},-2)$, if in addition $f\in L^{1}([0,T];L^{p}_{v}L^{1}_{x})$ for some $p>\frac{d}{d-|2+\gamma|}$, then we have
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U_{t}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}t\leq\sqrt{CC_{A}},$ (3.6)
where $C\lesssim\|\langle v\rangle^{|2+\gamma|}f\|_{L^{1}([0,T]\times\mathbb{R}^{2d})}+\|\langle v\rangle^{|2+\gamma|}f\|_{L^{1}([0,T];L^{p}_{v}L^{1}_{x})}$.
@@
We first show (3.5). By Cauchy–Schwarz inequality, we have
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U_{t}|\big(\langle v\rangle^{1+\frac{|\gamma|}{2}}+\langle x\rangle\big){\,\rm d}\eta{\,\rm d}t$ (3.7)
$\leq\Big(\int_{0}^{T}\int_{\mathbb{R}^{4d}}\frac{|U|^{2}}{ff_{*}\kappa}{\,\rm d}\eta{\,\rm d}t\Big)^{\frac{1}{2}}\times$
$\quad\Big(\int_{0}^{T}\int_{\mathbb{R}^{4d}}\big(\langle v\rangle^{2+|\gamma|}+\langle x\rangle^{2}\big){ff_{*}}\kappa{\,\rm d}\eta{\,\rm d}t\Big)^{\frac{1}{2}}$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U_{t}||v-v_{*}|^{1+\frac{\gamma}{2}}{\,\rm d}\eta{\,\rm d}r$
$\leq{C_{A}}^{\frac{1}{2}}\Big(\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa{ff_{*}}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t\Big)^{\frac{1}{2}}.$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}\kappa{ff_{*}}|v-v_{*}|^{2+\gamma}{\,\rm d}\eta{\,\rm d}t\lesssim 1+\|f\|_{L^{1}([0,T];L^{p}_{v}L^{1}_{x})}.$